OpenStudy (anonymous):

if (|x|-1)/(|x|-2)> or =0,,then find the solution set of x,where x is real number

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

work in cases if \(x>0\) this is \[\frac{x-1}{x-2}\geq0\] whereas if \(x<0\) this is \[\frac{-x-1}{-x-2}\geq 0\]

OpenStudy (blake57roger):

metal